### v2

# WeTok: Powerful Discrete Tokenization for High-Fidelity Visual Reconstruction

Abstract
Visual tokenizer is a critical component for vision generation. However, the existing tokenizers often face unsatisfactory trade-off between compression ratios and reconstruction fidelity. To fill this gap, we introduce a powerful and concise WeTok tokenizer, which surpasses the previous leading tokenizers via two core innovations. (1) Group-wise lookup-free Quantization (GQ). We partition the latent features into groups, and perform lookup-free quantization for each group. As a result, GQ can efficiently overcome memory and computation limitations of prior tokenizers, while achieving a reconstruction breakthrough with more scalable codebooks. (2) Generative Decoding (GD). Different from prior tokenizers, we introduce a generative decoder with a prior of extra noise variable. In this case, GD can probabilistically model the distribution of visual data conditioned on discrete tokens, allowing WeTok to reconstruct visual details, especially at high compression ratios. Extensive experiments on mainstream benchmarks show superior performance of our WeTok. On the ImageNet 50k validation set, WeTok achieves a record-low zero-shot rFID (WeTok: $\mathbf{0.12}$ vs. FLUX-VAE: $\mathbf{0.18}$ vs. SD-VAE 3.5: $\mathbf{0.19}$) with a $\mathbf{400\%}$ compression ratio. Furthermore, our highest compression model achieves a zero-shot rFID of $3.49$ with a compression ratio of 768, outperforming Cosmos (384) $4.57$ which has only 50% compression rate of ours. Code and models are available: https://github.com/zhuangshaobin/WeTok.
Figure 1: Zero-shot reconstruction comparison with state-of-the-art tokenizers. (a) Our WeTok establishes a new state-of-the-art trade-off between compression and reconstruction performance among the compared methods. (b) WeTok achieves a significant improvement in reconstruction quality over previous discrete tokenizers such as VQ-VAE and Open-MAGVIT2.
1 Introduction
In visual generation, the high computational cost of pixel-based data is a central challenge (Chen et al., 2020b; Rombach et al., 2022b). Visual tokenizers are a key solution that uses an encoder to compress an image into a compact latent representation and a decoder to reconstruct it (Kingma & Welling, 2013; Rezende et al., 2014), allowing generative models to operate efficiently in this latent space (Rombach et al., 2022a). These tokenizers are broadly divided into two categories: continuous (Kingma & Welling, 2013) and discrete (Van Den Oord et al., 2017). Continuous tokenizers map images to a continuous latent space, while discrete tokenizers employ a quantizer to produce a finite set of codes. This architectural difference introduces a critical trade-off. Discrete tokenizers can achieve a higher compression ratio, but this efficiency often comes at the cost of lower reconstruction fidelity compared to continuous methods. This leads to a natural question: Can we build a discrete tokenizer that can maintain high compression as well as achieve high-fidelity reconstruction?
To achieve this goal, two critical issues must be resolved. (1) Scalable Codebook. To minimize quantization error of discrete tokenizers, the existing methods attempt to enlarge the codebook (Yu et al., 2024a; Zhao et al., 2024b; Sun et al., 2024). In particular, the Lookup-Free Quantization (LFQ) (Yu et al., 2024a) quantizes the latent features directly, which largely increases the codebook size for better reconstruction. However, the substantial memory and computational overhead required to manage such a large codebook during training hinders further scalability. (2) Generative Modeling. Discrete tokenizers are inherently deterministic. Rather than modeling data distribution of images, decoder is trained to reconstruct the expected value of images (Esser et al., 2020), corresponding to the latent codes from encoder. Such a manner is limited to capture rich diversity and fine details in the original images, leading to unsatisfactory reconstruction, particularly at high compression ratios.
To fill the gap, we inrtdouce a powerful discrete tokenizer, namely WeTok, as shown in Fig. 1 (b). Specifically, we introduce two concise designs to solve the issues above. First, we develop a Group-Wise Lookup-Free Quantization (GQ), which partitions the latent feature of each image into groups, and performs lookup-free quantization for each group. As shown in Tab. 2, our GQ addresses the challenge in LFQ where entropy loss (Chang et al., 2022; Jansen et al., 2019) causes memory usage to grow with the codebook size, while yielding superior reconstruction performance. Second, we introduce a Generative Decoder (GD), which mimics the GAN-style generator by adding a prior of extra noise variable. As shown in Fig. 7, GD can effectively model data distribution of input images, allowing to reconstruct visual details at high compression ratios.
Finally, we conduct extensive experiments on mainstream benchmarks, via scaling WeTok across group size, model size, and training data size. Moreover, we pre-train our WeTok on a 400M general-domain dataset across multiple compression ratios. As illustrated in Fig. 1 (a), WeTok consistently outperforms the state-of-the-art continuous and discrete tokenizers with a $\mathbf{400\%}$ compression ratio, e.g., rFID on ImageNet 50k validation set: WeTok: $\mathbf{0.12}$ vs. FLUX-VAE: $\mathbf{0.18}$ (Batifol et al., 2025) vs. SD-VAE 3.5: $\mathbf{0.19}$ (Esser et al., 2024a). Furthermore, our highest compression model also achieves the superior reconstruction performance, e.g., rFID on ImageNet 50k validation set: WeTok: $3.59$ vs. Cosmos: $4.57$ (Agarwal et al., 2025), while Cosmos ($384$) only has 50% compression ratio of our WeTok ($768$), showing effectiveness and efficiency of WeTok.
2 Related Work
2.1 Continuous Tokenizer
Generative modeling in the pixel space typically requires extensive compute resources (Chen et al., 2020a; Ho et al., 2020). Subsequent works (Rombach et al., 2022b; Podell et al., 2023; Peebles & Xie, 2023; Esser et al., 2024b; Batifol et al., 2025; Zhuang et al., 2025; 2024; Esser et al., 2024a; Zha et al., 2025; Chen et al., 2025) adopt VAE (Kingma & Welling, 2013), which projects visual content from pixels to latent features, achieving efficient and photo-realistic visual generation at high resolution. FLUX-VAE (Batifol et al., 2025) shows the state-of-the-art performance in both reconstruction quality and generalization ability across all continuous tokenizers. However, continuous tokenizer is criticized for its low compression rate, because latent features are usually stored and calculated in float32 or bfloat16. Therefore, discrete tokenizers that can store data in int or bool seem to be more promising in terms of compression capabilities.
2.2 Discrete Tokenizer
Traditional works like VQVAE (Van Den Oord et al., 2017) and VQGAN (Esser et al., 2021) employ vector-quantization (VQ) to transform visual input into discrete tokens. But they both suffer from low reconstruction quality caused by quantization error and instability of the codebook utilization. To overcome these drawbacks, subsequent works have explored several paths. One line of work introduces specific optimization strategies or modules to improve performance (Lee et al., 2022b; Shi et al., 2024; Zhu et al., 2024; Yu et al., 2024b). Another line of work focuses on mitigating the training instability when scaling up the codebook size by using grouped codebooks (Ma et al., 2025; Jia et al., 2025; Zhang et al., 2025). These methods split the input feature into groups along the channel dimension, where each group is then looked up using a sub-codebook. However, despite these improvements, VQ-based tokenizers still introduce additional inference and training costs due to the lookup operation (Yu et al., 2021b; Lee et al., 2022b; Fang et al., 2025). MAGVIT-v2 (Yu et al., 2024a) introduces Lookup-Free Quantization (LFQ) to address this extra cost and proposes the entropy loss (Chang et al., 2022; Jansen et al., 2019) to ensure the utilization of the codebook. However, the entropy loss causes unaffordable memory cost as it scales linearly with the codebook size, limiting the further expansion of the codebook. Binary Spherical Quantization (BSQ) (Zhao et al., 2024a) is proposed to mitigate this memory issue by assuming independence between the bits of the binary code, while this strong assumption can lead to performance degradation compared to LFQ. Concurrent to the above work, the GQ in our WeTok does not rely on explicit codebooks, and can eliminate the memory usage caused by entropy loss while having better performance than LFQ.
2.3 Autoregressive Visual Generation
The autoregressive (AR) modeling paradigm, which underpins modern Large Language Models (LLMs) (Vaswani et al., 2017), has been successfully adapted for visual generation (Chen et al., 2020a), where models learn to predict sequences of discrete tokens for images (Ramesh et al., 2021b; Ding et al., 2021; Liu et al., 2024) and videos (Hong et al., 2022; Kondratyuk et al., 2023). Recent AR models (Sun et al., 2024; Team, 2024; Wu et al., 2025a; Wang et al., 2024b; Liu et al., 2025) achieve remarkable image quality, highlighting the significant potential of this paradigm. Notably, the success of AR models is critically dependent on the visual tokenizer. Therefore, we adopt our WeTok to the LlamaGen (Sun et al., 2024) framework to enable high-fidelity and efficient autoregressive generation. This shows that our WeTok is not only capable of compression, but its compressed features are also suitable for generative models.
3 Method
Figure 2: WeTok with Group-Wise Lookup-Free Quantization and Generative Decoder.
In this section, we first establish the necessary preliminaries for discrete tokenization. We then introduce the Group-Wise Lookup-Free Quantization (GQ) to unify Lookup-Free Quantization (LFQ) (Yu et al., 2024a) and Binary Spherical Quantization (BSQ) (Zhao et al., 2024b). Finally, we present a Generative Decoder (GD) specifically engineered for high-compression scenarios to reconstruct high-fidelity outputs from the compact representations generated by our GQ.
3.1 Preliminaries
Vector Quantized Variational Autoencoder (Esser et al., 2020). Given an image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, VQVAE first compresses it into latent feature $\mathcal{U}=\mathcal{E}(\mathcal{I})$, $\mathcal{U}\in\mathbb{R}^{h\times w\times d}$, through the encoder $\mathcal{E}$. Then it is quantized into latent codes $\mathcal{Q}$ by searching the nearest neighbor in the codebook $\mathcal{C}\in\mathbb{R}^{K\times d}=[\mathbf{c}_{1},\mathbf{c}_{2},\dots,\mathbf{c}_{K}]^{\top}$,
$\mathcal{Q}[i,j]=\mathop{\arg\min}_{\mathbf{c}\in{\mathcal{C}}}\|\mathcal{U}[i,j]-\mathbf{c}\|^{2},$ (1)
Notice that here we define $\mathcal{Q}$ to have a backward gradient only with respect to $\mathcal{C}$, and introduce $\mathcal{U}_{\mathcal{Q}}=\mathcal{U}+\text{sg}[\mathcal{Q}-\mathcal{U}]$ as a variable that shares the same value as $\mathcal{Q}$ but has a backward gradient only with respect to $\mathcal{U}$. Here, $\text{sg}[\cdot]$ denotes the stop-gradient operation.
Finally, the intermediate quantized result $\mathcal{Q}$ is reconstructed into image space $\hat{\mathcal{I}}=\mathcal{G}(\mathcal{U}_{\mathcal{Q}})$ through the decoder $\mathcal{G}$. The loss function of VQ-VAE consists of the following five parts,
$\mathcal{L}_{\text{VQVAE}}=\underbrace{\|\mathcal{I}-\hat{\mathcal{I}}\|^{2}}_{\text{Recon. Loss}}+\underbrace{\|\mathcal{Q}-\text{sg}[\mathcal{U}]\|^{2}}_{\text{Codebook Loss}}+\alpha\underbrace{\|\mathcal{U}-\text{sg}[\mathcal{Q}]\|^{2}}_{\text{Commitment Loss}}+\beta\underbrace{\mathcal{L}_{\text{LPIPS}}(\mathcal{I},\hat{\mathcal{I}})}_{\text{Perceptual Loss}}+\gamma\underbrace{\mathcal{L}_{\text{GAN}}(\mathcal{U}_{\mathcal{Q}})}_{\text{GAN Loss}},$ (2)
where perceptual loss (Zhang et al., 2018) and GAN loss are introduced for better visual quality.
Lookup-Free Quantization (Yu et al., 2024a). LFQ introduces an implicit and learning-free codebook $\mathcal{C}_{\text{LFQ}}=\{-1,1\}^{d}$ to perform lookup-free quantization on each channel of latent feature,
$\mathcal{Q}[i,j,k]=\text{sign}(\mathcal{U}[i,j,k]).$ (3)
Since the codebook in LFQ is fixed, there is no need for codebook loss during LFQ training. To address the issue of codebook utilization collapse in VQVAE, LFQ introduces the following entropy loss in replace of the commitment loss in equation 2,
$\mathcal{L}_{\text{Entropy}}(\mathcal{U})=\underbrace{\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}H\left(q(\mathbf{c}|\mathcal{U}[i,j])\right)}_{\text{Token Entropy Loss}}-\zeta\underbrace{H\left(\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}q(\mathbf{c}|\mathcal{U}[i,j])\right)}_{\text{Codebook Entropy Loss}}.$ (4)
where $H(X)=-\sum_{Xi\in X}X_{i}\text{log}X_{i}$ , $q(\mathbf{c}|\mathcal{U}[i,j])$ denote the conditional distribution of $\mathbf{c}$ given $\mathcal{U}[i,j]$ , and $\zeta$ refers to the weight of codebook entropy loss.
Binary Spherical Quantization (Zhao et al., 2024b). When increasing the codebook size, i.e., increasing $d$, the $H(\cdot)$ calculation in the $\mathcal{L}_{\text{Entropy}}$ of LFQ will lead to significant memory consumption. To alleviate this issue, BSQ seeks to handle each binary of LFQ seperately. BSQ first rewrites the token entropy loss into binary seperated form as $\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}\sum_{k=1}^{d}H(q_{B}(\mathbf{c}[k]|\mathcal{U}[i,j,k]))$. Next, BSQ rewrites the codebook entropy loss as $\sum_{k=1}^{d}H(q_{B}(\mathbf{c}[k]|\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}\mathcal{U}[i,j,k]))$ by assuming the approximation $\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}q(\mathbf{c}|\mathcal{U}[i,j]))\approx\prod_{k=1}^{d}\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}q_{B}(\mathbf{c}[k]|\mathcal{U}[i,j,k]))$. Both operations reduce the variable space of the $H(\cdot)$ calculation from $\{-1,1\}^{d}$ to the linear combination of $d$ $\{-1,1\}$, significantly decreasing memory consumption. However, as the binary approximation for the codebook entropy loss proposed by BSQ would introduce additional errors, the performance of BSQ is inferior to that of LFQ under the same codebook size.
3.2 Group-Wise Lookup-Free Quantization
As shown in Eq. 4, the computational cost of $H(\cdot)$ increases linearly with the codebook size. To simultaneously address it and the optimization error issue of BSQ, we propose grouping the latent features first and then performing lookup-free quantization. As shown in Fig. 2, we group the latent features from the channel dimension, reshape $\mathcal{U}$ into $\mathcal{U}_{\text{G}}\in\mathbb{R}^{h\times w\times g\times d^{\prime}}$, where $d=gd^{\prime}$ and $g$ is a newly added dimension representing the number of groups and $d^{\prime}$ is group channel. For $k$-th group, there is a non-learnable grouped codebook $\mathcal{C}_{\text{GQ},k}=\{-1,1\}^{d^{\prime}}$. The conditional probability after grouping can be formulated as
$q(\mathbf{c}|\mathcal{U}[i,j])=\prod_{k=1}^{g}q_{\text{G}}(\mathbf{c}_{k}|\mathcal{U}_{G}[i,j,k]),$ (5)
where $\mathbf{c}_{k}$ refers to the $k$-th latent code after $\mathbf{c}$ is divided into $g$ parts, i.e., $\mathbf{c}_{k}=\mathbf{c}[(k-1)d^{\prime}+1:kd^{\prime}]$. Considering the additivity of entropy and Eq. 5, we can rewrite the token entropy loss term as
$\mathcal{L}_{\text{Token Entropy Loss}}=\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}H(q(\mathbf{c}|\mathcal{U}[i,j]))=\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}\sum_{k=1}^{g}H(q_{\text{G}}(\mathbf{c}_{k}|\mathcal{U_{\text{G}}}[i,j,k])).$ (6)
We transform the original token entropy loss calculation from the $\{-1,1\}^{d}$ space into grouped token entropy loss which is a linear combination of entropy from $g$ $\{-1,1\}^{d^{\prime}}$ spaces. This change eliminates the token entropy loss as the memory bottleneck.
However, for the codebook entropy loss, the $H(\sum\cdot)$ operation prevents us from leveraging the additivity of entropy to decompose the $\{-1,1\}^{d}$ space into a linear combination of multiple subspaces. We propose the assumption that
$\sum_{i=1}^{h}\sum_{j=1}^{w}q(\mathbf{c}|\mathcal{U}[i,j])=\sum_{i=1}^{h}\sum_{j=1}^{w}\prod_{k=1}^{g}q_{\text{G}}(\mathbf{c}_{k}|\mathcal{U}_{\text{G}}[i,j,k])\approx\prod_{k=1}^{g}\sum_{i=1}^{h}\sum_{j=1}^{w}q_{\text{G}}(\mathbf{c}_{k}|\mathcal{U}_{\text{G}}[i,j,k]).$ (7)
Thus, we leverage the additivity of entropy to transform the codebook entropy loss into
$\mathcal{L}_{\text{Codebook Entropy Loss}}=H(\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}q(\mathbf{c}|\mathcal{U}[i,j]))=\sum_{k=1}^{g}H(\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}q_{\text{G}}(\mathbf{c}_{k}|\mathcal{U}_{\text{G}}[i,j,k])).$ (8)
Similar to the token entropy loss, we transform the calculation of the codebook entropy loss from the $\{-1,1\}^{d}$ space into a grouped codebook entropy loss, which becomes a linear combination of entropy from $g$ $\{-1,1\}^{d^{\prime}}$ spaces.
Governed by the group number $g$, GQ provides a tunable trade-off between approximation accuracy and memory cost. When $g=1$, GQ degenerates to the LFQ formulation, which uses no approximation and thus achieves high precision at the cost of substantial memory consumption. Conversely, when $g=d$, GQ mirrors the BSQ approach, which employs a strong approximation to minimize memory usage but consequently suffers from significant optimization error. Experiments in Sec. 4.2 demonstrate that by selecting an appropriate intermediate value for $g$, GQ can significantly reduce memory overhead while introducing only minimal optimization error. This allows GQ to surpass even the standard LFQ baseline. Furthermore, this tunable design provides the flexibility to effectively scale the codebook to a virtually unlimited size.
3.3 Generative Decoder
Unlike the continuous tokenizer, the decoders in previous discrete (Esser et al., 2021; Yu et al., 2024a) tokenizers fit deterministic transformations. Although GAN loss is employed during training as shown in Eq. 2 and its specific form is as follows
$\mathcal{L}_{\text{GAN}}(\mathcal{U}_{\mathcal{Q}})=\text{log}(1-\mathcal{D}(\mathcal{G}(\mathcal{U}_{\mathcal{Q}}))),$ (9)
where $\mathcal{D}$ is the discriminator. However, the GAN loss only serves to assist in improving the perceptual quality. In scenarios with high compression rates, the correspondence between $\mathcal{U}_{\mathcal{Q}}$ and the ground truth is likely not unique. In such cases, we need the decoder to become a generative model. As shown in Fig. 2, we randomly sample $\mathbf{z}\in\mathcal{N}(\mathbf{0},\mathbf{I})$, concatenate it with $\mathcal{U}_{\mathcal{Q}}$ along the channel dimension, and then feed the result into the decoder. In this way, the GAN loss is subsequently reformulated as
$\mathcal{L}_{\text{GAN}}(\mathcal{U}_{\mathcal{Q}})=\mathbb{E}_{\mathbf{z}\in\mathcal{N}(\mathbf{0},\mathbf{I})}[\text{log}(1-\mathcal{D}(\mathcal{G}(\mathbf{z},\mathcal{U}_{\mathcal{Q}})))],$ (10)
where $\mathcal{U}_{\mathcal{Q}}$ serves as a condition for the generation process. The development from Eq. 9 to Eq. 10 is not merely concatenating $\mathbf{z}$ to the input of the decoder. More importantly, the learning objective of the decoder shifts to modeling the transformation from Gaussian noise conditioned on $\mathcal{U}_{\mathcal{Q}}$ to the ground truth distribution, turning it into a GD.
Figure 3: Quantization method ablation. GQ and LFQ are significantly better than BSQ.
Figure 4: Number of group ablation. $G$ refers to the number of group. The reconstruction performance of the model increases significantly with the increase of $G$.
Figure 5: Model architecture ablation. $C$ and $B$ refer to the number of base channel and residual block respectively. $C=256$ and $B=4$ achieve the best reconstruction performance.
Figure 6: Training data and learning rate schedule ablation. GD refers to general-domain data. The model trained on general-domain data is not as good as the model trained on in-distribution data in terms of distribution fitting metrics, but has better generalization on PSNR and SSIM. The effect of consistent learning schedule is significant compared with warm up + cosine decay.
To ensure training stability, we employ a two-stage training process. In the first stage, we train our WeTok with the reconstruction loss, i.e., Eq. 2, 6 and 8. In the second stage, we adapt the model for generative tasks. Specifically, we expand the channel dimension of the conv_in layer in decoder to accept z as additional input. To preserve the powerful reconstruction capabilities learned in the first stage, the newly added weights for these channels are zero-initialized. This strategy ensures that at the beginning of the second stage, the decoder’s behavior is identical to its pre-trained state. Consequently, this versatile approach allows our GD to be effectively integrated with various pre-trained discrete tokenizers that share a similar architectural foundation.
4 Experiments
Figure 7: Qualitative ablation of GD on MS-COCO val2017. The images reconstructed by the model with GD are obviously more fidelity and natural.
Figure 8: Parameter ablation experiments of the Autoregressive model based on WeTok. Compared with Open-MAGVIT2, the performance of the AR model based on WeTok improves more significantly with the increase in the number of parameters.
Table 1: Memory usage ablation. We set $d^{\prime}$==88 in GFQ. OOM refers to out of memory.
Method $d=8$ $d=16$ $d=24$ $d=32$ $d=40$
LFQ 10.5 GB 10.6 GB OOM OOM OOM
BSQ 10.5 GB 10.5 GB 10.6 GB 10.6 GB 10.6 GB
GFQ 10.5 GB 10.6 GB 10.6 GB 10.6 GB 10.6 GB
Table 2: Generative decoder modeling ablation. Stage2 refers to generative decoder modeling.
Stage1 Stage2 rFID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
5.37 0.17 0.54 20.53
3.90 0.16 0.55 20.72
Table 3: Reconstruction evaluation on $\mathbf{256\times 256}$ ImageNet 50K validation set. All models are trained on ImageNet. WeTok achieves SOTA results on different downsampling rates.
[…]
VQGAN (Esser et al., 2020) 2D $16\times 16$ 16 $256\times 256$ $1024$ 8.30 19.51 -
VQGAN (Esser et al., 2020) 2D $16\times 16$ 16 $256\times 256$ $16384$ 4.99 20.00 -
SD-VQGAN (Rombach et al., 2022b) 2D $16\times 16$ 16 $256\times 256$ $16384$ 5.15 - -
MaskGIT (Chang et al., 2022) 2D $16\times 16$ 16 $256\times 256$ $1024$ 2.28 - -
ReVQ (Zhang et al., 2025) 2D $16\times 16$ 16 $256\times 256$ $65536$ 2.57 21.69 -
LlamaGen (Sun et al., 2024) 2D $16\times 16$ 16 $256\times 256$ $32768$ 2.26 20.59 85%
LlamaGen (Sun et al., 2024) 2D $16\times 16$ 16 $256\times 256$ $16384$ 2.19 20.79 97%
ReVQ (Zhang et al., 2025) 2D $16\times 16$ 16 $256\times 256$ $262144$ 2.05 21.96 -
VAR (Tian et al., 2024) 2D $16\times 16$ 16 $256\times 256$ $4096$ - 21.30 97%
IBQ (Shi et al., 2025) 2D $16\times 16$ 16 $256\times 256$ $16384$ 1.37 22.35 96%
Open-MAGVIT2 (Luo et al., 2024) 2D $16\times 16$ 16 $256\times 256$ $2^{18}$ 1.17 22.64 100%
IBQ (Shi et al., 2025) 2D $16\times 16$ 16 $256\times 256$ $262144$ 1.00 20.30 84%
VFMTok (Zheng et al., 2025) 1D $256$ - $256\times 256$ $16384$ 0.89 - 100%
GigaTok (Xiong et al., 2025) 1D $256$ - $256\times 256$ $16384$ 0.79 21.65 -
AliTok (Wu et al., 2025c) 1D $273$ - $256\times 256$ $4096$ 0.84 - -
MGVQ (Jia et al., 2025) 2D $16\times 16$ 16 $256\times 256$ $2^{52}$ 0.64 23.71 100%
WeTok (Ours) 2D $16\times 16$ 16 $256\times 256$ $2^{32}$ 0.61 24.50 100%
ViT-VQGAN (Yu et al., 2021a) 2D $32\times 32$ 8 $256\times 256$ $8192$ 1.28 - -
OmiTokenizer-VQ (Wang et al., 2024a) 2D $32\times 32$ 8 $256\times 256$ $8192$ 1.11 - -
LlamaGen (Sun et al., 2024) 2D $32\times 32$ 8 $256\times 256$ $16384$ 0.59 24.45 -
Open-MAGVIT2 (Luo et al., 2024) 2D $32\times 32$ 8 $\times 128 2182^{18}$ 0.34 27.02 100%
BSQ (Zhao et al., 2024a) 1D $1024$ - $256\times 256$ $2^{18}$ 1.14 25.36 100%
BSQ (Zhao et al., 2024a) 1D 1024 - $256\times 256$ $2^{36}$ 0.45 28.14 100%
MGVQ (Jia et al., 2025) 2D $32\times 32$ 8 $256\times 256$ $2^{52}$ 0.31 28.42 100%
WeTok (Ours) 2D $32\times 32$ 8 $256\times 256$ $2^{32}$ 0.19 29.69 100%
Table 4: Zero-shot reconstruction performance on ImageNet 50k validation set and MS-COCO val2017. The tokenizers are trained with large scale general-domain datasets. Text in gray signifies the results directly from Cosmos (Agarwal et al., 2025) report. Our WeTok achieves the best performance on both resolution settings.
Method Tokenizer Training Ratio Compression MS-COCO 2017 Imagenet-1k
[…]
Resize $256\times 256$
DALL-E dVAE (Ramesh et al., 2021a) Discrete 103M 18 118 48.60 26.97 0.08 32.63 27.31 0.79
Cosmos (Agarwal et al., 2025) Discrete - 16 384 11.97 19.22 0.48 4.57 19.93 0.49
Show-o (Xie et al., 2024) Discrete 35M 16 473 9.26 20.90 0.59 3.50 21.34 0.59
Open-MAGVIT2-I-PT (Luo et al., 2024) Discrete 100M 16 439 7.93 22.21 0.62 2.55 22.21 0.62
LlamaGen (Sun et al., 2024) Discrete 70M 16 439 8.40 20.28 0.55 2.47 20.65 0.54
Open-MAGVIT2-I-PT (Luo et al., 2024) Discrete 100M 16 341 6.76 22.31 0.65 1.67 22.70 0.64
WeTok (Ours) Discrete 400M 32 768 8.94 20.31 0.55 3.49 20.77 0.55
BSQ (Zhao et al., 2024b) Discrete 1B - 219 - - - 3.81 24.12 0.66
QLIP-B (Zhao et al., 2025) Discrete 1B - 219 - - - 3.21 23.16 0.63
QLIP-L (Zhao et al., 2025) Discrete 1B - 168 - - - 1.46 25.36 0.69
SD-VAE 1.x (Rombach et al., 2022a) Continuous 1B 8 24 5.94 23.21 0.69 1.22 23.54 0.68
SD-VAE 1.x (Rombach et al., 2022a) Discrete 1B 8 110 5.75 24.17 0.70 1.13 24.48 0.69
WeTok (Ours) Discrete 400M 16 192 4.41 24.44 0.74 0.60 24.77 0.73
QLIP-B (Zhao et al., 2025) Discrete 1B - 55 - - - 0.70 26.79 0.79
SD-VAE 2.x (Rombach et al., 2022a) Continuous 6B 8 24 4.26 26.62 0.77 0.70 26.90 0.76
SDXL-VAE (Podell et al., 2023) Continuous >6B 8 24 3.93 27.08 0.80 0.67 27.37 0.78
UniTok (Ma et al., 2025) Discrete 1B 16 64 - - - 0.41 - -
WeTok (Ours) Discrete 400M 8 48 2.18 29.49 0.89 0.20 29.63 0.88
SD-VAE 3.5 (Esser et al., 2024b) Continuous - 8 6 1.66 31.08 0.90 0.19 31.19 0.90
FLUX-VAE (Labs, 2024) Continuous - 8 6 1.35 32.32 0.93 0.18 32.74 0.92
WeTok (Ours) Discrete 400M 8 24 1.43 32.00 0.93 0.12 32.06 0.93
Original Resolution
DALL-E dVAE (Ramesh et al., 2021a) Discrete 103M 18 118 55.07 25.15 0.75 36.84 25.46 0.74
Cosmos (Agarwal et al., 2025) Discrete - 16 384 7.51 20.45 0.52 1.93 20.56 0.51
Cosmos (Agarwal et al., 2025) Discrete - 16 384 7.23 20.45 0.53 2.52 20.49 0.52
Open-MAGVIT2-I-PT (Luo et al., 2024) Discrete 100M 16 439 6.65 21.61 0.57 1.39 21.74 0.56
Open-MAGVIT2-I-PT (Luo et al., 2024) Discrete 100M 16 341 5.10 22.18 0.60 0.78 22.24 0.59
WeTok (Ours) Discrete 400M 32 768 7.83 20.47 0.52 2.03 20.54 0.51
SD-VAE 1.x (Rombach et al., 2022a) Continuous 1B 8 24 5.94 21.68 0.64 1.35 21.99 0.63
SD-VAE 1.x (Rombach et al., 2022a) Discrete 1B 8 110 6.07 22.54 0.65 1.23 22.82 0.64
WeTok (Ours) Discrete 400M 16 192 3.80 23.70 0.67 0.40 23.75 0.67
SD-VAE 2.x (Rombach et al., 2022a) Continuous 6B 8 24 4.63 24.82 0.72 0.78 25.08 0.71
SDXL-VAE (Podell et al., 2023) Continuous >6B 8 24 4.23 25.11 0.74 0.72 25.38 0.73
WeTok (Ours) Discrete 400M 8 48 2.09 27.50 0.82 0.18 27.54 0.82
SD-VAE 3.5 (Esser et al., 2024b) Continuous - 8 6 1.64 28.35 0.86 0.24 28.39 0.86
WeTok (Ours) Discrete 400M 8 24 1.46 29.47 0.88 0.12 29.51 0.88
4.1 Implementation
Datasets. To thoroughly evaluate our WeTok, we perform large-scale training on two datasets: (i) 1.2M ImageNet (Russakovsky et al., 2014) training set; (ii) 400M general-domain dataset. To make state-of-the-art comparison, we then evaluate WeTok performance on the ImageNet 50k validation set and MS-COCO 2017 validation set (Lin et al., 2014b). This multi-benchmark evaluation ensures a robust assessment of our model’s capabilities across different domains and data distributions. Unless otherwise stated, we conduct a series of ablation studies on the ImageNet training set. Besides, we train the class-to-image model on the ImageNet training set and test it on the validation set.
Settings. WeTok adopts the architecture proposed in Open-MAGVIT2 (Luo et al., 2024), employing the CNN architecture for encoder, decoder, and discriminator. Images are randomly cropped to $256\times 256$ for training. For the ablation study, all models are trained for 250K steps using a consistent set of hyperparameters. Including a fixed learning rate of 1e-4, the most global batch size is 128, and the Adam (Kingma & Ba, 2014) optimizer with $\beta_{1}=0.5$ and $\beta_{2}=0.9$. For large-scale training, we used the same Adam optimizer settings, but increased the most global batch size to 1024. In this phase, other hyperparameters were individually tuned for each model to achieve optimal performance. For class-to-image generation, we adopt the generative transformer architecture in LlamaGen (Sun et al., 2024). All experiments were carried out using the H20 GPU with Pytorch. More details in Sup. A.
4.2 Ablation Study
We conducted a comprehensive ablation study to validate the key components of WeTok. We first verified the effectiveness of our two important algorithms, GQ and GD. Then we ablate the performance improvements brought to WeTok by various dimensions, including the number of groups $g$ in GQ, model architecture, training data, and learning rate schedule. In addition, we conduct parameter ablation experiments on the AR model based on WeTok. More implementation details in Sup. A.1. We employ validation loss, rFID (Heusel et al., 2017), LPIPS (Zhang et al., 2018), SSIM (Wang et al., 2004), and PSNR to evaluate the quality of ablation study.
Quantization method. We conduct an ablation study of quantization methods under the same compression ratio, $i.e.$, $d=gd^{\prime}=16$. As shown in Tab. 2, our GQ not only has almost no increase in GPU memory usage due to quantization like BSQ. In Fig. 3, we set the same compression ratio for 3 different quantization methods (LFQ: gg==11, $d^{\prime}$==$16$; BSQ: gg==$16$, $d^{\prime}$==11; GQ: gg==22, $d^{\prime}$==88), GQ performs better than LFQ and far exceeds BSQ.
Figure 9: Qualitative comparison of $\mathbf{512\times 512}$ image reconstruction on TokBench.
Figure 10: WeTok-AR-XL generated samples at $\mathbf{256\times 256}$ resolution.
Table 5: Class-conditional generation on $\mathbf{256\times 256}$ ImageNet. * specifies the generated images are $384\times 384$ and are resized to 256×256 for evaluation.
[…]
Diffusion ADM (Dhariwal & Nichol, 2021) 554M 10.94 101.0 0.69 0.63
CDM (Ho et al., 2021) - 4.88 158.7 - -
LDM-4 (Rombach et al., 2022b) 400M 3.60 247.7 - -
DiT-XL/2 (Peebles & Xie, 2022) 675M 2.27 278.2 0.83 0.57
AR VQGAN (Esser et al., 2020) 227M 18.65 80.4 0.78 0.26
VQGAN (Esser et al., 2020) 1.4B 15.78 74.3 - -
VQGAN-re (Esser et al., 2020) 1.4B 5.20 280.3 - -
ViT-VQGAN (Yu et al., 2021a) 1.7B 4.17 175.1 - -
ViT-VQGAN-re (Yu et al., 2021a) 1.7B 3.04 227.4 - -
RQTran. (Lee et al., 2022a) 3.8B 7.55 134.0 - -
RQTran.-re (Lee et al., 2022a) 3.8B 3.80 323.7 - -
VAR VAR-d16 (Tian et al., 2024) 310M 3.30 274.4 0.84 0.51
VAR-d20 (Tian et al., 2024) 600M 2.57 302.6 0.83 0.56
VAR-d24 (Tian et al., 2024) 1.0B 2.09 312.9 0.82 0.59
VAR-d30 (Tian et al., 2024) 2.0B 1.92 323.1 0.82 0.59
AR LlamaGen-L (Sun et al., 2024) 343M 3.07 256.06 0.83 0.52
LlamaGen-XL (Sun et al., 2024) 775M 2.62 244.08 0.80 0.57
LlamaGen-XXL (Sun et al., 2024) 1.4B 2.34 253.90 0.80 0.59
LlamaGen-L (Sun et al., 2024) 343M 3.80 248.28 0.83 0.51
LlamaGen-XL (Sun et al., 2024) 775M 3.39 227.08 0.81 0.54
LlamaGen-XXL (Sun et al., 2024) 1.4B 3.09 253.61 0.83 0.53
UniTok (Ma et al., 2025) 1.4B 2.51 216.7 0.82 0.57
Open-MAGVIT2-AR-B (Luo et al., 2024) 343M 3.08 258.26 0.85 0.51
Open-MAGVIT2-AR-L (Luo et al., 2024) 804M 2.51 271.70 0.84 0.54
Open-MAGVIT2-AR-XL (Luo et al., 2024) 1.5B 2.33 271.77 0.84 0.54
WeTok-AR-XL (Ours) 1.5B 2.31 276.55 0.84 0.55
Generative decoder. When the performance of the model is saturated after the first stage of reconstruction training, we conduct the second stage of generative training. As shown in Tab. 2, the results show that GD can continue to improve the reconstruction performance of the model, especially in rFID. In addition, we present qualitative ablation results in Fig. 7. After converting the decoder to a generative model, the reconstructed images are more realistic and consistent with visual logic, demonstrating the effectiveness of GD.
Number of group in GQ. We increase the number of groups $g$ by a power of 2. As shown in Fig. 4, the results show that as $g$ increases, the reconstruction performance of the model continues to increase significantly and does not encounter the memory bottleneck like LFQ.
Model architecture. We performed ablation studies to scale up our discrete tokenizer, focusing on the number of base channels and residual blocks in the encoder and decoder. As shown in Fig. 5, across 9 different settings, a configuration with 256 base channels and 4 residual blocks achieves the best reconstruction performance. The encoder and decoder for this optimal architecture contain 198M and 261M parameters, respectively.
Training data. We compare the model trained on the 1.2M ImageNet training set and the other trained on a larger 400M general-domain dataset. As shown in Fig. 6, the model trained on the general-domain dataset achieved a better validation loss and SSIM and PSNR scores. However, it yielded worse rFID and LPIPS scores. We attribute it to the data distribution gap. The ImageNet training and validation sets are in-distribution, whereas the general-domain data are significantly more diverse. This reveals the trade-off between generalization and performance on specific in-distribution evaluation metrics.
Learning rate schedule. A learning rate schedule composed of a warm-up phase followed by a cosine decay is widely adopted in training models. However, we found that this convention may be suboptimal for the training of discrete tokenizers. As illustrated in Fig. 6, the model trained with a constant learning rate demonstrates markedly better performance. Based on this, we exclusively adopted the constant learning rate schedule for all subsequent large-scale training in WeTok.
Parameter of autoregressive model. As shown in Fig. 8, We conduct ablation experiments on the parameter size of the WeTok-based AR model and compared it with an Open-MAGVIT2-based AR model. The experimental results show that the performance of the WeTok-based AR model is slightly inferior to that of the Open-MAGVIT2-based AR model when the parameter size is small. However, as the parameter size increases, the WeTok-based AR model surpasses the Open-MAGVIT2-based AR model in all indicators.
4.3 Comparison with state-of-the-art
Visual Reconstruction. We first evaluate WeTok’s performance in an in-distribution setting, where both the training and validation data originate from the ImageNet dataset. This establishes our model’s core effectiveness before addressing the more complex challenge of distribution gaps. As shown in Tab. 3, WeTok achieves state-of-the-art reconstruction performance, outperforming existing methods across different downsampling ratios. Subsequently, we compare WeTok against other methods, with all models being trained on a large-scale general-domain dataset. To ensure a fair and unbiased comparison, we have carefully filtered this dataset to exclude any data samples in ImageNet and MS-COCO Lin et al. (2014a). This crucial step prevents potential data leakage and ensures that the evaluation accurately measures the models’ ability to generalize from a broad data distribution to the specific target domain. As shown in Tab. 4, our WeTok shows the state-of-the-art reconstruction performance in a wide range of compression ratio scenarios. It not only shows the strongest performance among discrete tokenizers, but also even surpasses the current strongest continuous tokenizers, FLUX-VAE (Labs, 2024) and SD-VAE 3.5 (Esser et al., 2024a). In addition, we present qualitative comparison results on the TokBench (Wu et al., 2025b) in Fig. 9. Compared to the widely used SDXL-VAE (Podell et al., 2023) and SD-VAE 1.5 (Rombach et al., 2022b), WeTok’s reconstruction performance is significantly superior under the same compression ratio.
Visual Generation. To evaluate WeTok’s capabilities beyond image reconstruction, we extend its application to visual generation. We adopt an autoregressive model modified from LlamaGen as Open-MAGVIT2 for this task, primarily due to the well-established scalability of such models. Specifically, we employ the in-distribution WeTok with $16\times$ downsampling in Tab. 3 as the image tokenizer. More training details can be found in Sup. A.2. As shown in Tab. 5, our WeTok-based generative model achieves state-of-the-art performance on the ImageNet 50K validation set. This result demonstrates that our WeTok is a highly effective tokenizer not only for image reconstruction but also for high-fidelity visual generation tasks. As shown in Fig. 10, we show the realistic and diverse image generation results of our WeTok-AR-XL.
5 Conclusion
In this paper, we introduce WeTok, a powerful discrete visual tokenizer designed to resolve the long-standing conflict between compression efficiency and reconstruction fidelity. Our primary contributions are the Group-Wise Lookup-Free Quantization (GQ) method, which provides a scalable and memory-efficient solution for codebooks, and Generative Decoder (GD) that excels at producing high-fidelity images even from highly compressed representations. Through extensive experiments, we demonstrated that WeTok consistently outperforms existing state-of-the-art discrete and continuous tokenizers in both in-distribution and zero-shot reconstruction tasks across a wide range of compression ratios. Furthermore, by integrating WeTok into an autoregressive framework, we achieved state-of-the-art performance in class-conditional image generation, confirming that its learned tokens are highly effective for downstream generative tasks. WeTok establishes a new performance standard, proving that discrete tokenizers can achieve superior reconstruction quality without compromising their inherent advantage in compression.
Appendix A More Implementation Details
A.1 Ablation Study
Quantization method. As shown in Tab. 8, 8 and 8.
Table 6: GFQ training setting.
config GFQ
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 2
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $250$
channel_mult [1,1,2,2,4]
channel 128
num_res_blocks 4
generative decoder False
per GPU batchsize 16
global batchsize 128
GPU number 8 H20
Table 7: LFQ training setting.
config LFQ
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 1
$d^{\prime}$ (group channel) 16
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $250$
channel_mult [1,1,2,2,4]
channel 128
num_res_blocks 4
generative decoder False
per GPU batchsize 16
global batchsize 128
GPU number 8 H20
Table 8: BSQ training setting.
config BSQ
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 16
$d^{\prime}$ (group channel) 1
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $250$
channel_mult [1,1,2,2,4]
channel 128
num_res_blocks 4
generative decoder False
per GPU batchsize 16
global batchsize 128
GPU number 8 H20
Generative decoder. As shown in Tab. 10 and 10.
Table 9: Stage-1 training setting.
config Stage-1
  training data general domain dataset
image size [256, 256]
data augmentation random crop
downsample $32\times 32$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $550$
channel_mult [1,1,2,2,4,8]
channel 256
num_res_blocks 4
generative decoder False
per GPU batchsize 6
global batchsize 1056
GPU number 176 H20
Table 10: Stage-2 training setting.
config Stage-2
  training data general domain dataset
image size [256, 256]
data augmentation random crop
downsample $32\times 32$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $550$
channel_mult [1,1,2,2,4,8]
channel 256
num_res_blocks 4
generative decoder True
per GPU batchsize 6
global batchsize 1056
GPU number 176 H20
Number of group in GQ. As shown in Tab. 16, 16, 16, 16, 16 and 16.
Table 11: 1 group training setting.
config 1 group
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 1
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $250$
channel_mult [1,1,2,2,4]
channel 128
num_res_blocks 4
generative decoder False
per GPU batchsize 16
global batchsize 128
GPU number 8 H20
Table 12: 2 group training setting.
config 2 group
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 2
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $250$
channel_mult [1,1,2,2,4]
channel 128
num_res_blocks 4
generative decoder False
per GPU batchsize 16
global batchsize 128
GPU number 8 H20
Table 13: 2 group training setting.
config 4 group
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $250$
channel_mult [1,1,2,2,4]
channel 128
num_res_blocks 4
generative decoder False
per GPU batchsize 16
global batchsize 128
GPU number 8 H20
Table 14: 8 group training setting.
config 8 group
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 8
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $250$
channel_mult [1,1,2,2,4]
channel 128
num_res_blocks 4
generative decoder False
per GPU batchsize 16
global batchsize 128
GPU number 8 H20
Table 15: 16 group training setting.
config 16 group
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 16
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $250$
channel_mult [1,1,2,2,4]
channel 128
num_res_blocks 4
generative decoder False
per GPU batchsize 16
global batchsize 128
GPU number 8 H20
Table 16: 32 group training setting.
config 32 group
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 32
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $250$
channel_mult [1,1,2,2,4]
channel 128
num_res_blocks 4
generative decoder False
per GPU batchsize 16
global batchsize 128
GPU number 8 H20
Model architecture. As shown in Tab. 25, 25, 25, 25, 25, 25, 25, 25, and 25.
Table 17: 128 channel 4 block training setting.
config 128 channel 4 block
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $250$
channel_mult [1,1,2,2,4]
channel 128
num_res_blocks 4
generative decoder False
per GPU batchsize 16
global batchsize 128
GPU number 8 H20
Table 18: 192 channel 4 block training setting.
config 192 channel 4 block
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $250$
channel_mult [1,1,2,2,4]
channel 192
num_res_blocks 4
generative decoder False
per GPU batchsize 8
global batchsize 128
GPU number 16 H20
Table 19: 256 channel 4 block training setting.
config 256 channel 4 block
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $250$
channel_mult [1,1,2,2,4]
channel 256
num_res_blocks 4
generative decoder False
per GPU batchsize 4
global batchsize 128
GPU number 32 H20
Table 20: 128 channel 8 block training setting.
config 128 channel 8 block
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $250$
channel_mult [1,1,2,2,4]
channel 128
num_res_blocks 8
generative decoder False
per GPU batchsize 8
global batchsize 128
GPU number 16 H20
Table 21: 128 channel 16 block training setting.
config 128 channel 16 block
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $250$
channel_mult [1,1,2,2,4]
channel 128
num_res_blocks 16
generative decoder False
per GPU batchsize 4
global batchsize 128
GPU number 32 H20
Table 22: 192 channel 8 block training setting.
config 192 channel 8 block
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $250$
channel_mult [1,1,2,2,4]
channel 192
num_res_blocks 8
generative decoder False
per GPU batchsize 4
global batchsize 128
GPU number 32 H20
Table 23: 256 channel 2 block training setting.
config 256 channel 2 block
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $250$
channel_mult [1,1,2,2,4]
channel 256
num_res_blocks 2
generative decoder False
per GPU batchsize 8
global batchsize 128
GPU number 16 H20
Table 24: 384 channel 4 block training setting.
config 384 channel 4 block
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $250$
channel_mult [1,1,2,2,4]
channel 384
num_res_blocks 4
generative decoder False
per GPU batchsize 2
global batchsize 128
GPU number 64 H20
Table 25: 256 channel 8 block training setting.
config 256 channel 8 block
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $250$
channel_mult [1,1,2,2,4]
channel 256
num_res_blocks 8
generative decoder False
per GPU batchsize 2
global batchsize 128
GPU number 64 H20
Training data. As shown in Tab. 27 and 27.
Table 26: ImageNet training setting.
config 1.2M
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $550$
channel_mult [1,1,2,2,4]
channel 128
num_res_blocks 4
generative decoder False
per GPU batchsize 16
global batchsize 128
GPU number 8 H20
Table 27: General-domain training setting.
config 400M
  training data general domain dataset
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $550$
channel_mult [1,1,2,2,4]
channel 128
num_res_blocks 4
generative decoder False
per GPU batchsize 16
global batchsize 128
GPU number 8 H20
Learning rate schedule. As shown in Tab. 28.
Table 28: Warm up + cosine decay learning rate schedule training setting.
config warm up + cosine decay
  training data general domain dataset
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule warm up + cosine decay
learning rate 1e-4
warmup steps 10000
cos decay end ratio 0.01
total steps $550$
channel_mult [1,1,2,2,4]
channel 128
num_res_blocks 4
generative decoder False
per GPU batchsize 16
global batchsize 128
GPU number 8 H20
Parameter of autoregressive model. As shown in Tab. 31, 31 and 31.
Table 29: LlamaGen Base training setting.
config LlamaGen Base
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema False
optimizer AdamW
optimizer momentum $\beta_{1},\beta_{2}{=}0.9,0.95$
weight decay 5e-2
learning rate schedule warm up + linear decay
learning rate 3e-4
warmup epochs 6
linear decay end ratio 0.1
total epochs 1000
dim 1024
num_head 16
trans_layers 24
cond_dim 1024
factorized_layers 2
factorized_kk 4
token_drop 0.1
residual_drop 0.1
per GPU batchsize 64
global batchsize 3072
GPU number 48 H20
Table 30: LlamaGen Large training setting.
config LlamaGen Large
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema False
optimizer AdamW
optimizer momentum $\beta_{1},\beta_{2}{=}0.9,0.95$
weight decay 5e-2
learning rate schedule warm up + linear decay
learning rate 3e-4
warmup epochs 6
linear decay end ratio 0.1
total epochs 1000
dim 1280
num_head 20
trans_layers 36
cond_dim 1280
factorized_layers 3
factorized_kk 4
token_drop 0.1
residual_drop 0.1
per GPU batchsize 32
global batchsize 3072
GPU number 96 H20
Table 31: LlamaGen X-Large training setting.
config LlamaGen X-Large
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema False
optimizer AdamW
optimizer momentum $\beta_{1},\beta_{2}{=}0.9,0.95$
weight decay 5e-2
learning rate schedule warm up + linear decay
learning rate 3e-4
warmup epochs 6
linear decay end ratio 0.1
total epochs 1000
dim 1536
num_head 24
trans_layers 48
cond_dim 1536
factorized_layers 4
factorized_kk 4
token_drop 0.1
residual_drop 0.1
per GPU batchsize 16
global batchsize 3072
GPU number 192 H20
A.2 Comparison with state-of-the-art
Visual Reconstruction. The settings of in-distribution comparison are shown in Tab. 33 and 33. The settings of general-domain comparison are shown in Tab. 37, 37, 37 and 37, .
Visual Generation. As shown in Tab. 31.
Table 32: Large scale training on ImageNet training set at 16 downsample ratio.
config IN-1K $16\times$ SOTA
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $400$
channel_mult [1,1,2,2,4]
channel 256
num_res_blocks 4
generative decoder False
per GPU batchsize 8
global batchsize 1024
GPU number 128 H20
Table 33: Large scale training on ImageNet training set at 8 downsample ratio.
config IN-1K $8\times$ SOTA
  training data IN-1K training set
image size [256, 256]
data augmentation random crop
downsample $8\times 8$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate 1e-4
warmup steps 0
cos decay end ratio 1
total steps $350$
channel_mult [1,2,2,4]
channel 256
num_res_blocks 4
generative decoder False
per GPU batchsize 8
global batchsize 1024
GPU number 128 H20
Table 34: Large scale training on general-domain dataset at 768 compression ratio.
config $768$ compression ratio SOTA
  training data general domain dataset
image size [256, 256]
data augmentation random crop
downsample $32\times 32$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate
1e-4 1e-4 1e-5
warmup steps 0
cos decay end ratio 1
total steps
$550$ $550$ $330$
channel_mult [1,1,2,2,4,8]
channel 256
num_res_blocks 4
generative decoder
False True True
per GPU batchsize 6
global batchsize 1056
GPU number 176 H20
Table 35: Large scale training on general-domain dataset at 192 compression ratio.
config $192$ compression ratio SOTA
  training data general domain dataset
image size [256, 256]
data augmentation random crop
downsample $16\times 16$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate
1e-4 1e-5 1e-6
warmup steps 0
cos decay end ratio 1
total steps
$470$ 90090 10010
channel_mult [1,1,2,2,4]
channel 256
num_res_blocks 4
generative decoder
False False False
per GPU batchsize 8
global batchsize 1024
GPU number 128 H20
Table 36: Large scale training on general-domain dataset at 48 compression ratio.
config $48$ compression ratio SOTA
  training data general domain dataset
image size [256, 256]
data augmentation random crop
downsample $8\times 8$
ema True
$g$ (group number) 4
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate
1e-4 1e-5 1e-6
warmup steps 0
cos decay end ratio 1
total steps
$200$ $200$ 20020
channel_mult [1,2,2,4]
channel 256
num_res_blocks 4
generative decoder False
per GPU batchsize 8
global batchsize 1024
GPU number 128 H20
Table 37: Large scale training on general-domain dataset at 24 compression ratio.
config $24$ compression ratio SOTA
  training data general domain dataset
image size [256, 256]
data augmentation random crop
downsample $8\times 8$
ema True
$g$ (group number) 8
$d^{\prime}$ (group channel) 8
optimizer Adam
optimizer momentum $\beta_{1},\beta_{2}{=}0.5,0.9$
weight decay 0
learning rate schedule consistent
learning rate
1e-4 1e-5 1e-5
warmup steps 0
cos decay end ratio 1
total steps
$300$ 50050 $400$
channel_mult [1,2,2,4]
channel 256
num_res_blocks 4
generative decoder False
per GPU batchsize 8
global batchsize
1024 1024 4096
GPU number
128 H20 128 H20 512 H20